Solve the following system of IVP: [11 x = Ax where A=[1 1 LO 0 0 0 -10 and x(0) = 4 [2.
Answer #1

x= An where A 2 o and xo J 0 2 Eigenvalues A are of 1 A-NI- ED O o 30 -N -l O 1-2 0 1 x o o -1-2 (-1-1) ( (1-4)2-1] 13-1 1-x)2x + 0 Oxi + 2, 26 - Тус - гр. 2 - 2 ) + м со -За ъб и, об 1 ? - . - 2(0) эб Os in Choosing My = ! Hena R = (oo 1). eigenvect.1 - -3 esthetik- ocru-lh -303 so ni=12 1350 sel (o, Mm, 23) 11,1,0) solution is (choosing My 3 M2 <) Now as = (A) u Gelit utOn adding first two equations Cty +(-(2+) = 2t4 263 = 6 (3=3 (2 + (go2 (2+3=2 62--1 Pulting in * (t) 1 2*7* 69(4)3200 + #